Experimental Volumes

Mixture

Volume (mL) Fe(NO3)3

Assigned Volume (mL) KSCN
Read Instructions Above

Volume (mL) KSCN
Read Instructions Above

Volume (mL) Water

You (Blank)

5.00

0.00

0.00

5.00

You (Volume 1)

5.00

0.5

Read Instructions Above!!

0.7

Try 1: ✔ 4.30

1 1

You (Volume 2)

5.00

2.2

2.2

Try 1: ✔ 2.80

1 1

You (Volume 2 Chilled)

5.00

2.2

2.2

2.80

Partner (Blank)

5.00

0.00

0.00

5.00

Partner (Volume 1)

5.00

3.2

3.9

Try 1: ✔ 1.10

1 1

Partner (Volume 2)

5.00

4.5

4.8

Try 1: ✔ 0.20

1 1
Spectrometer Readings

Mixture

%T (1st Reading)

%T (2nd Reading)

Average %T

Absorbance

You (Blank)

96.9

98.7

Try 1: ✔ 97.8

1 1

0.00

You (Volume 1)

34.7

34.5

Try 1: ✔ 34.6

1 1
NG 1
You (Volume 2)

29.2

29.4

NG 1
NG 1
You (Volume 2 Chilled)

53.2

42.7

NG 1
NG 1

Partner (Blank)

92.6

92.1

NG 1

0.00

Partner (Volume 1)

83.5

83.7

NG 1
NG 1

Partner (Volume 2)

56.7

57.5

NG 1
NG

1

How do i get absorbance ?
Answer #1

We know that

%T = I/Io * 100

we know  T = I/Io

so, T = %T / 100 .............(1)

and also we know

A = -log T .................(2)

Now putting the value of T from equation (1) in equation (2), we get;

A = -log (%T / 100) .......................(3)

Now putting the value of %T in equation (3), we get

for blank we have

A = -log(97.8/100)

= log(100/97.8)

= 0.00

for volume 1, we have

A = -log(34.6/100)

= log(100/34.6) = 0.46

for volume 2, we have %T average = (29.2+29.4)/2 =29.3

A = -log(29.3/100)

= log(100/29.3) = 0.53

similarly we can calculate Absorbance for each %T reading.

hope this helps!

NOTE: A = absrobance

T = transmittance

%T = %transmittance

I = intensity of the light transmitted

Io = intensity of the light entering